### v1

@@ 2.3.2 Distribution of $\mathcal{T}^{\ell}_{k}$
In this part, we describe the distribution of the uniform $k$-point subtree $\mathcal{T}^{\ell}_{k}$. We recall Aldous’ coalescent construction and then reformulate it in terms of a piecewise Kingman representation. The latter formulation will play a key role in Section 5.1. In particular, it makes explicit that the finite-dimensional convergence of the tree sequence is equivalent to the convergence of the associated coalescent dynamics to Kingman’s coalescent.
Aldous’ coalescent construction
In [1998BROWNIAN], the uniform $k$-point subtree $\mathcal{T}^{\ell}_{k}$ is described via a coalescent process of $k$ particles. Fix a continuous profile $\ell$ with $h(\ell)<\infty$ and $I(\ell)=1$. Think of “time” $t$ running downward from $\infty$ to 0. Each of $k$ particles is born at independent random times with probability density $\ell(x)\,\mathrm{d}x$. As time decreases, clusters of particles merge according to the following rule: in the interval $t,t-\mathrm{d}t]$, each pair of existing clusters coalesces into one cluster with probability $\frac{4}{\ell(t)}\,\mathrm{d}t$. At time $t=0$, all remaining clusters merge into a single cluster.
@@
To obtain an ordered tree, we assign a left-right ordering at each branching point. Whenever two clusters merge at height $t>0$, one is designated as the left subtree and the other as the right subtree, with the choice made uniformly at random. At time $t=0$, when the final merge occurs, a uniformly random ordering is again assigned. This procedure yields a rooted ordered real tree. Aldous [1998BROWNIAN] showed that the law of this tree coincides with the law of the uniform $k$-point subtree $\mathcal{T}^{\ell}_{k}$.
For a general continuous profile $\ell$ with $h(\ell)<\infty$, one easily checks that the law of $\mathcal{T}^{\ell}_{k}$ is given by the same coalescent construction, except that the birth density $\ell(x)\,\mathrm{d}x$ is replaced by its normalized form $\frac{\ell(x)}{I(\ell)}\,\mathrm{d}x$.
@@
For real numbers $x_{1}>x_{2}>\cdots>x_{k}>0$, the piecewise Kingman’s coalescent $\big(Z^{(k)}_{s}=Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s\geq 0\big)$ with rate $\frac{4}{\ell(s)}$ is defined as:
@@
  • Recursively, for each $i=1,\ldots,k-1$, conditionally on $\big(Z^{(k)}_{s}:s>x_{i}\big)$, the process $Z^{(k)}_{s}$ on the interval $s\in(x_{i+1},x_{i}]$ evolves as a Kingman’s coalescent started with $\lim_{s\downarrow x_{i}}Z^{(k)}_{s}+1$ lineages, where each pair of lineages coalesces at rate $\frac{4}{\ell(s)}$. Here the height parameter $s$ is interpreted as backward time, so the coalescent runs as $s$ decreases.
@@
For any $k\geq 1$ and real numbers $x_{1}>x_{2}>\cdots>x_{k}>0$, conditionally on the leaf heights being $x_{1},\ldots,x_{k}$, the process $Y^{(k)}_{s}$ evolves as the piecewise Kingman’s coalescent $\big(Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s\geq 0\big)$.
@@
    Sample a piecewise Kingman’s coalescent process $\big(Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s\geq 0\big)$.
@@
    Conditional on $\{Z^{(k)}_{s}\}_{s\geq 0}$, sample uniformly from the set of rooted ordered real trees whose lineage counts at each height agree with $Z^{(k)}$.
@@ 2.4 Distribution of uniform $k$-point subtrees induced by $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$
Proof 2.17.
By Proposition 2.10, the total time duration of $W^{\ell^{\sigma}}$ equals $I(\ell^{\sigma})$. Using (8), we have a.s.
@@
Since $\sigma>0$ on $(0,h(\ell))$, it follows that $\sigma(W^{\ell^{\sigma}}_{s})^{2}>0$ for a.e. $s$, hence $\alpha$ is a.s. strictly increasing.
@@
The fact that $W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}\in C_{\mathrm{bridge}}$ then follows from Proposition 2.10 (ii).
@@
    Sample a piecewise Kingman coalescent $\big(Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s\geq 0\big)$ in which each pair of lineages coalesces at rate $\frac{4}{\ell^{\sigma}(s)}$.
@@
    Conditional on $\{Z^{(k)}_{s}\}_{s\geq 0}$, sample uniformly from the set of rooted ordered real trees whose lineage counts at each height agree with $Z^{(k)}$.
@@
$W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}=W^{\ell^{\sigma}}_{I(\ell^{\sigma})\,\phi_{t}},\qquad\phi_{t}:=\frac{\alpha^{-1}(I(\ell)t)}{I(\ell^{\sigma})}.$
@@
$\mathcal{T}^{\ell,\sigma}(t_{1},\ldots,t_{k})\;\cong\;\mathcal{T}^{\ell^{\sigma}}(\phi_{t_{1}},\ldots,\phi_{t_{k}}).$ (9)
@@
$\mathbb{E}\big[f(\mathcal{T}^{\ell,\sigma}_{k})\big]=\left(\frac{I(\ell^{\sigma})}{I(\ell)}\right)^{k}\mathbb{E}\!\big[f(\mathcal{T}^{\ell^{\sigma}}_{k})\,\sigma(\mathcal{T}^{\ell^{\sigma}}_{k})^{2}\big],$ (10)
@@
Since
$\phi^{-1}(s)=\frac{\alpha(I(\ell^{\sigma})s)}{I(\ell)},\qquad\mathrm{d}\phi^{-1}(s)=\frac{I(\ell^{\sigma})}{I(\ell)}\,\sigma(W^{\ell^{\sigma}}_{I(\ell^{\sigma})s})^{2}\,\mathrm{d}s,$
substituting into (11) yields
$\mathbb{E}\big[f(\mathcal{T}^{\ell,\sigma}_{k})\big]$ $=\left(\frac{I(\ell^{\sigma})}{I(\ell)}\right)^{k}\int_{0}^{1}\cdots\int_{0}^{1}\mathbb{E}\big[f(\mathcal{T}^{\ell^{\sigma}}(s_{1},\ldots,s_{k}))\big]\prod_{j=1}^{k}\sigma(W^{\ell^{\sigma}}_{I(\ell^{\sigma})s_{j}})^{2}\,\mathrm{d}s_{1}\cdots\,\mathrm{d}s_{k}$
$=\left(\frac{I(\ell^{\sigma})}{I(\ell)}\right)^{k}\mathbb{E}\big[f(\mathcal{T}^{\ell^{\sigma}}_{k})\sigma(\mathcal{T}^{\ell^{\sigma}}_{k})^{2}\big].$
This completes the proof.
Proof 2.23 (Proof of Theorem 2.20).
@@
$\mathbb{E}\big[f(\mathring{\mathcal{T}}^{\ell,\sigma}_{k})\big]=\left(\frac{I(\ell^{\sigma})}{I(\ell)}\right)^{k}\,\mathbb{E}\!\big[f(\mathcal{T}^{\ell^{\sigma}}_{k})\,\sigma(\mathcal{T}^{\ell^{\sigma}}_{k})^{2}\big].$
Together with (10), this implies $\mathring{\mathcal{T}}^{\ell,\sigma}_{k}\overset{d}{=}\mathcal{T}^{\ell,\sigma}_{k}$, completing the proof.
@@ 3 Coalescent process
In this section, we first derive preparatory moment and probability estimates for elementary coalescence events (Section 3.1); these bounds will be used repeatedly in Sections 35. We then introduce the coalescent process arising in Cannings trees, show that it forms a Markov chain, and prove a coming-down-from-infinity property that serves as a central ingredient for the tightness of $\widetilde{\mathscr{T}}^{n}$.
We begin with some notation. Denote by $x_{1}^{n,s},\dots,x_{q_{n}(s)}^{n,s}$ the vertices at generation $s$ of $\mathscr{T}^{n}$, listed according to the lexicographic order. Let $(u_{1}^{n,s},\dots,u_{q_{n}(s)}^{n,s})$ be a uniform random permutation of $(x_{1}^{n,s},\dots,x_{q_{n}(s)}^{n,s})$, independent across different generations and independent of $\mathscr{T}^{n}$.
For any vertex $x$, we write $\xi(x)$ for the number of children of $x$ and $\mathfrak{p}(x)$ for the parent of $x$. In particular, we have $\nu_{i}^{n,s}=\xi(x_{i}^{n,s})$.
@@ 3.1 Coalescent probability
In this subsection, we introduce coalescent events in Cannings trees and express their probabilities in terms of the moments of the offspring counts $\nu$. We then estimate these moments, which in turn yields quantitative estimates for the coalescent probabilities.
@@
$M^{s}=M^{n,s}:=\Big\{f:A\to\{1,\ldots,q_{n}(s)\}\,:\,A\subset\{1,\ldots,q_{n}(s+1)\}\Big\},$ (12)
@@
$\text{Coal}^{s}(f)=\text{Coal}^{n,s}(f):=\Big\{\forall\,j\in\mathrm{Dom}(f),\ \mathfrak{p}(u^{n,s+1}_{j})=x^{n,s}_{f(j)}\Big\}.$ (13)
Suppose the range of $f$ is $\{j_{1}<j_{2}<\cdots<j_{m}\}\subset\{1,\ldots,q_{n}(s+1)\}$. We then define
@@ 3.1.1 Linking coalescent probability to offspring moments
${\mathbb{P}}_{n}\big(\text{Coal}^{s}(f)\big)=\frac{1}{\big(q_{n}(s+1)\big)_{\#\mathrm{Dom}(f)}}\,\mathbb{E}_{n}\!\Bigg[\prod_{i=1}^{m}\big(\nu^{s}_{j_{i}}\big)_{N_{i}}\Bigg],$ (14)
@@
$\frac{\prod_{i=1}^{m}(v^{s}_{j_{i}})_{N_{i}}}{(q_{n}(s+1))_{\#\mathrm{Dom}(f)}}.$
Taking expectation over $(\nu^{s}_{j_{i}})$ yields
${\mathbb{P}}_{n}\big(\text{Coal}^{s}(f)\big)=\frac{1}{(q_{n}(s+1))_{\#\mathrm{Dom}(f)}}\,{\mathbb{E}}_{n}\Bigg[\prod_{i=1}^{m}\big(\nu^{s}_{j_{i}}\big)_{N_{i}}\Bigg],$
as claimed.
@@
For $n\geq 1$, $1\leq s<h(q_{n})$, $1\leq l\leq m\leq q_{n}(s+1)$, it holds:
  1. (1)
    ${\mathbb{P}}_{n}\!\Big(\{\mathfrak{p}(u_{1}^{s+1})=\cdots=\mathfrak{p}(u_{l}^{s+1})\}\cap\{\mathfrak{p}(u_{l}^{s+1}),\dots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\}\Big)$
    $=$ $\frac{(q_{n}(s))_{m-l+1}}{(q_{n}(s+1))_{m}}\,{\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})_{l}\prod_{i=2}^{m-l+1}\nu_{i}^{s}\Big],$
  2. (2)
    ${\mathbb{P}}_{n}\!\Big(\mathfrak{p}(u_{1}^{s+1})=\mathfrak{p}(u_{2}^{s+1})\neq\mathfrak{p}(u_{3}^{s+1})=\mathfrak{p}(u_{4}^{s+1})\Big)=\frac{(q_{n}(s))_{2}}{(q_{n}(s+1))_{4}}\,{\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})_{2}(\nu_{2}^{s})_{2}\Big].$
For later use in the proof of the main theorem, when applying part (1) of Corollary 3.3, we will mainly need the cases $l=1$ or $2$. Following the same argument as in the proof of Proposition 3.1, Corollary 3.3 admits the following refinement.
@@
For $n,K\geq 1$, $1\leq s<h(q_{n})$ and $1\leq l\leq q_{n}(s+1)$, it holds:
  • (1)
    ${\mathbb{P}}_{n}\!\Big(\mathfrak{p}(u_{1}^{s+1})=\cdots=\mathfrak{p}(u_{l}^{s+1}),\;\xi(\mathfrak{p}(u_{1}^{s+1}))\leq K\Big)=\frac{q_{n}(s)}{(q_{n}(s+1))_{l}}\,{\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})_{l}\mathbf{1}\{\nu_{1}^{s}\leq K\}\Big],$
  • (2)

### v2

@@ 2.3.2 Distribution of $\mathcal{T}^{\ell}_{k}$
In this part, we describe the distribution of the uniform $k$-point subtree $\mathcal{T}^{\ell}_{k}$. We recall Aldous’ construction and then reformulate it in terms of a piecewise Kingman’s representation. The latter formulation will play a key role in Section 5.1. In particular, it makes explicit that the finite-dimensional convergence of the tree sequence is equivalent to the convergence of the associated coalescent dynamics to Kingman’s coalescent.
Aldous’ construction via a coalescent process
In [3], the uniform $k$-point subtree $\mathcal{T}^{\ell}_{k}$ is described via a coalescent process of $k$ particles. Fix a continuous profile $\ell$ with $h(\ell)<\infty$ and $I(\ell)=1$. Think of “time” $t$ running downward from $\infty$ to 0. Each of $k$ particles is born at independent random times with probability density $\ell(x)\,\mathrm{d}x$. As time decreases, clusters of particles merge according to the following rule: in the interval $t,t-\mathrm{d}t]$, each pair of existing clusters coalesces into one cluster with probability $\frac{4}{\ell(t)}\,\mathrm{d}t$. At time $t=0$, all remaining clusters merge into a single cluster.
@@
To obtain an ordered tree, we assign a left-right ordering at each branching point. Whenever two clusters merge at height $t>0$, one is designated as the left subtree and the other as the right subtree, with the choice made uniformly at random. At time $t=0$, when the final merge occurs, a uniformly random ordering is again assigned. This procedure yields a rooted ordered real tree. Aldous [3] showed that the law of this tree coincides with the law of the uniform $k$-point subtree $\mathcal{T}^{\ell}_{k}$.
For a general continuous profile $\ell$ with $h(\ell)<\infty$, one easily checks that the law of $\mathcal{T}^{\ell}_{k}$ (recall Definition 2.11) is given by the same coalescent construction, except that the birth density $\ell(x)\,\mathrm{d}x$ is replaced by its normalized form $\frac{\ell(x)}{I(\ell)}\,\mathrm{d}x$.
@@
For real numbers $x_{1}>x_{2}>\cdots>x_{k}>0$, the piecewise Kingman’s coalescent $\big(Z^{(k)}_{s}=Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s>0\big)$ with rate $\frac{4}{\ell(s)}$ is defined as:
@@
  • Recursively, for each $i=1,\ldots,k$, conditionally on $\big(Z^{(k)}_{s}:s>x_{i}\big)$, the process $Z^{(k)}_{s}$ on the interval $s\in(x_{i+1},x_{i}]$ (with $x_{k+1}:=0$) evolves as a Kingman’s coalescent started with $\lim_{s\downarrow x_{i}}Z^{(k)}_{s}+1$ lineages, where each pair of lineages coalesces at rate $\frac{4}{\ell(s)}$. Here the height parameter $s$ is interpreted as backward time, so the coalescent runs as $s$ decreases.
@@
For any $k\geq 1$ and real numbers $x_{1}>x_{2}>\cdots>x_{k}>0$, conditionally on the leaf heights being $x_{1},\ldots,x_{k}$, the process $Y^{(k)}_{s}$ evolves as the piecewise Kingman’s coalescent $\big(Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s>0\big)$.
@@
    Sample a piecewise Kingman’s coalescent process $\big(Z^{(k)}_{s}=Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s\geq 0\big)$.
@@
    Conditional on $(Z^{(k)}_{s})_{s>0}$, sample uniformly from the set of rooted ordered real trees whose lineage counts at each height agree with $Z^{(k)}$.
@@ 2.4 Distribution of uniform $k$-point subtrees induced by $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$
Proof.
By Proposition 2.10, the total time duration of $W^{\ell^{\sigma}}$ equals $I(\ell^{\sigma})$. Using (2.5), we have a.s.
@@
Since $\sigma>0$ on $(0,h(\ell))$, it follows that $\sigma(W^{\ell^{\sigma}}_{s})^{2}>0$ for a.e. $s$, hence $\alpha$ is a.s. strictly increasing. ∎
@@
The fact that $W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}\in C_{\mathrm{bridge}}$ then follows from Proposition 2.10 (ii). ∎
@@
    Sample a piecewise Kingman coalescent $Z^{(k)}=\big(Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s>0\big)$ in which each pair of lineages coalesces at rate $\frac{4}{\ell^{\sigma}(s)}$.
@@
    Conditionally on $Z^{(k)}$, sample uniformly from the set of rooted ordered real trees whose lineage counts at each height agree with $Z^{(k)}$.
@@
$W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}=W^{\ell^{\sigma}}_{I(\ell^{\sigma})\phi(t)},\qquad\phi(t):=\frac{\alpha^{-1}(I(\ell)t)}{I(\ell^{\sigma})}.$
@@
$\mathcal{T}^{\ell,\sigma}(t_{1},\ldots,t_{k})\text{ is isomorphic to }\mathcal{T}^{\ell^{\sigma}}\big(\phi({t_{1}}),\ldots,\phi({t_{k}})\big).$ (2.6)
@@
$\mathbb{E}\big[f(\mathcal{T}^{\ell,\sigma}_{k})\big]=\left(\frac{I(\ell^{\sigma})}{4I(\ell)}\right)^{k}\mathbb{E}\!\big[f(\mathcal{T}^{\ell^{\sigma}}_{k})\,\sigma(\mathcal{T}^{\ell^{\sigma}}_{k})^{2}\big],$ (2.7)
@@
By the definition of $\phi$,
$\phi^{-1}(s)=\frac{\alpha(I(\ell^{\sigma})s)}{I(\ell)},\qquad\mathrm{d}\phi^{-1}(s)=\frac{I(\ell^{\sigma})}{4I(\ell)}\,\sigma(W^{\ell^{\sigma}}_{I(\ell^{\sigma})s})^{2}\,\mathrm{d}s,$
substituting into (2.8) yields
$\mathbb{E}\big[f(\mathcal{T}^{\ell,\sigma}_{k})\big]$ $=\left(\frac{I(\ell^{\sigma})}{4I(\ell)}\right)^{k}\int_{0}^{1}\cdots\int_{0}^{1}\mathbb{E}\big[f(\mathcal{T}^{\ell^{\sigma}}(s_{1},\ldots,s_{k}))\big]\prod_{j=1}^{k}\sigma(W^{\ell^{\sigma}}_{I(\ell^{\sigma})s_{j}})^{2}\,\mathrm{d}s_{1}\cdots\,\mathrm{d}s_{k}$
$=\left(\frac{I(\ell^{\sigma})}{4I(\ell)}\right)^{k}\mathbb{E}\big[f(\mathcal{T}^{\ell^{\sigma}}_{k})\sigma(\mathcal{T}^{\ell^{\sigma}}_{k})^{2}\big].$
This completes the proof. ∎
Proof of Theorem 2.18.
@@
$\mathbb{E}\big[f(\mathring{\mathcal{T}}^{\ell,\sigma}_{k})\big]=\left(\frac{I(\ell^{\sigma})}{4I(\ell)}\right)^{k}\,\mathbb{E}\!\big[f(\mathcal{T}^{\ell^{\sigma}}_{k})\,\sigma(\mathcal{T}^{\ell^{\sigma}}_{k})^{2}\big].$
Together with (2.7), this implies $\mathring{\mathcal{T}}^{\ell,\sigma}_{k}\overset{d}{=}\mathcal{T}^{\ell,\sigma}_{k}$, thus completing the proof. ∎
@@ 3 Coalescent process
In this section, we first derive preparatory moment and probability estimates for elementary coalescence events (Section 3.1); these bounds will be used repeatedly in Sections 35. We then introduce the coalescent process arising in Cannings trees, show that it forms a Markov chain, and prove a coming-down-from-infinity (CDFI) property that serves as a central ingredient for the tightness of $\widetilde{\mathscr{T}}^{n}$.
We begin with some notation. Denote by $x_{1}^{n,s},\dots,x_{q_{n}(s)}^{n,s}$ the vertices at generation $s$ of $\mathscr{T}^{n}$, listed according to the lexicographic order. Let
[…]
@@ 3.1 Coalescent probability
In this subsection, we introduce coalescent events in Cannings trees and express their probabilities in terms of moments of the offspring counts $\nu$. We then estimate these moments, which in turn yield quantitative estimates for the coalescent probabilities.
@@
$M^{s}=M^{n,s}:=\Big\{f:A\to\{1,\ldots,q_{n}(s)\}\,:\,A\subset\{1,\ldots,q_{n}(s+1)\}\Big\},$ (3.1)
@@
$\text{Coal}^{s}(f)=\text{Coal}^{n,s}(f):=\Big\{\forall\,j\in\mathrm{Dom}(f),\ \mathfrak{p}(u^{n,s+1}_{j})=x^{n,s}_{f(j)}\Big\}.$ (3.2)
Suppose the range of $f$ is $\{j_{1}<j_{2}<\cdots<j_{m}\}\subset\{1,\ldots,q_{n}(s)\}$. We then define
@@ 3.1.1 Linking coalescent probability to offspring moments
${\mathbb{P}}_{n}\big(\text{Coal}^{s}(f)\big)=\frac{1}{\big(q_{n}(s+1)\big)_{\#\mathrm{Dom}(f)}}\,\mathbb{E}_{n}\!\Bigg[\prod_{i=1}^{m}\big(\nu^{s}_{j_{i}}\big)_{N_{i}}\Bigg],$ (3.3)
@@
$\frac{\prod_{i=1}^{m}(\nu^{s}_{j_{i}})_{N_{i}}}{(q_{n}(s+1))_{\#\mathrm{Dom}(f)}}.$
Taking expectation over $(\nu^{s}_{j_{i}})$ yields (3.3). ∎
@@
For $n\geq 1$, $1\leq s<h(q_{n})$, $1\leq l\leq m\leq q_{n}(s+1)$, it holds that
${\mathbb{P}}_{n}\!\Big(\{\mathfrak{p}(u_{1}^{s+1})=\cdots=\mathfrak{p}(u_{l}^{s+1})\}\cap\{\mathfrak{p}(u_{l}^{s+1}),\dots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\}\Big)=\,\frac{(q_{n}(s))_{m-l+1}}{(q_{n}(s+1))_{m}}\,{\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})_{l}\prod_{i=2}^{m-l+1}\nu_{i}^{s}\Big],$ (3.4)
${\mathbb{P}}_{n}\!\Big(\mathfrak{p}(u_{1}^{s+1})=\mathfrak{p}(u_{2}^{s+1})\neq\mathfrak{p}(u_{3}^{s+1})=\mathfrak{p}(u_{4}^{s+1})\Big)=\frac{(q_{n}(s))_{2}}{(q_{n}(s+1))_{4}}\,{\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})_{2}(\nu_{2}^{s})_{2}\Big].$ (3.5)
For later use in the proof of the main theorem, when applying (3.4), we will mainly need the cases $l=1$ or $2$. Following the same argument as in the proof of Proposition 3.1, Corollary 3.2 admits the following refinement.
@@
For $n,K\geq 1$, $1\leq s<h(q_{n})$ and $1\leq l\leq q_{n}(s+1)$, it holds that:
${\mathbb{P}}_{n}\!\Big(\mathfrak{p}(u_{1}^{s+1})=\cdots=\mathfrak{p}(u_{l}^{s+1}),\;\xi(\mathfrak{p}(u_{1}^{s+1}))\leq K\Big)=\frac{q_{n}(s)}{(q_{n}(s+1))_{l}}\,{\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})_{l}\,\mathbf{1}\{\nu_{1}^{s}\leq K\}\Big];$